how do you solve
x+y+z=-54
x=-6y
z=14y

Answer by checkley77(12844)

x+y+z=-54
x=-6y
z=14y
Replace x with (-6y) & replace z with (14y) in the first equation & solve for y.
-6y+y+14y=-54
9y=-54
y=-54/9
y=-6 ans.
x=-6*-6=36 ans.
z=14*-6=-84 ans.
Proof:
36-6+14*-6=-54
36-6-84=-54
36-90=-54
-54=-54